### v1

# DemoRank: Selecting Effective Demonstrations for Large Language Models in Ranking Task

Abstract
Recently, there has been increasing interest in applying large language models (LLMs) as zero-shot passage rankers. However, few studies have explored how to select appropriate in-context demonstrations for the passage ranking task, which is the focus of this paper. Previous studies mainly apply a demonstration retriever to retrieve demonstrations and use top-$k$ demonstrations for in-context learning (ICL). Although effective, this approach overlooks the dependencies between demonstrations, leading to inferior performance of few-shot ICL in the passage ranking task. In this paper, we formulate the demonstration selection as a retrieve-then-rerank process and introduce the DemoRank framework. In this framework, we first use LLM feedback to train a demonstration retriever and construct a novel dependency-aware training samples to train a demonstration reranker to improve few-shot ICL. The construction of such training samples not only considers demonstration dependencies but also performs in an efficient way. Extensive experiments demonstrate DemoRank’s effectiveness in in-domain scenarios and strong generalization to out-of-domain scenarios. Our codes are available at https://github.com/8421BCD/DemoRank.
DemoRank: Selecting Effective Demonstrations for Large Language Models in Ranking Task
Wenhan Liu, Yutao Zhu,  and Zhicheng Dou Gaoling School of Artificial Intelligence, Renmin University of China lwh@ruc.edu.cn, yutaozhu94@gmail.com, dou@ruc.edu.cn
1 Introduction
Large language models (LLM) have demonstrated remarkable performance across a spectrum of natural language processing (NLP) tasks. Recently, there has been significant interest in using LLMs for passage ranking tasks Zhuang et al. (2023a); Sun et al. (2023); Qin et al. (2023). A typical approach is relevance generation, which judges the relevance of a query-passage pair in a pointwise manner. This method prompts LLMs to assess the relevance of a passage to a query by generating responses such as “Yes” or “No”. The relevance score is then computed based on the log-likelihood of these responses. This approach has been demonstrated to be effective in previous studies Zhuang et al. (2023a); Liang et al. (2022).
Figure 1: Compared with choosing top-2 demonstrations ($z_{1}$ and $z_{2}$), the combination of $z_{1}$ and $z_{5}$ provides richer and more diverse query-passage relationships, thus yielding better relevance assessment.
In-context learning (ICL) has been proved as an emergent ability of LLMs Wei et al. (2022), enabling them to adapt to specific tasks through several task demonstrations (i.e., input-output examples). Many studies have investigated the optimal selection of demonstrations for NLP tasks Lu et al. (2022); Zhang et al. (2022); Li et al. (2023); Wang et al. (2023); Xu et al. (2024), highlighting the importance of tailored demonstrations in achieving high performance. However, the application of ICL to passage ranking tasks has not been extensively studied. Given the complex nature of passage ranking, ICL presents a challenging yet promising opportunity to enhance LLMs’ performance. Consequently, this study aims to develop effective demonstration selection strategies to optimize the application of ICL in passage ranking.
A widely-used and effective approach for demonstration selection is training a demonstration retriever using LLM’ feedback Wang et al. (2023); Rubin et al. (2022); Li et al. (2023); Cheng et al. (2023); Scarlatos and Lan (2023); Luo et al. (2023). This approach first utilizes an LLM to score some demonstration candidates based on LLM’s likelihood of producing the correct output given each candidate and the input, and choose positive and negative candidates based on scores for retriever training. Following this technique line, we propose to train a demonstration retriever based on LLM’s feedback tailored for passage ranking task.
In the inference stage, a common practice Wang et al. (2023) is to use the trained retriever to obtain a list of demonstrations and concatenate the top-retrieved ones together in the prompt for ICL. Despite its effectiveness in NLP tasks, directly extending it into the passage ranking task may result in sub-optimal performance. The main challenge lies in the complex nature of the query-passage relationship in passage ranking, which may require a combination of multiple demonstrations to provide effective information for understanding such a relationship. Figure 1 shows an example of such a problem. When selecting 2-shot demonstrations for the current input (a relevant query-passage pair), existing methods Wang et al. (2023); Rubin et al. (2022) will choose the top-2 demonstrations ($z_{1}$ and $z_{2}$) returned by the retriever. However, we deem that combining $z_{1}$ and $z_{5}$ is more suitable for this case. This is because $z_{1}$ and $z_{5}$ have more distinct queries and opposite outputs (relevance label), which provide LLM with richer and more diverse query-passage relationship signals, thus contributing more to the relevance assessment. This example shows the insufficiency of pure relevance-based demonstration selection in the few-shot LLM-based passage ranking task. In this paper, we transform the problem of selecting the optimal $k$-shot demonstrations from initially retrieved $n$ demonstrations into a demonstration ranking problem and propose to use LLM’s feedback to train a novel dependency-aware demonstration reranker, making the top-ranked ones more suitable in the few-shot ICL for passage ranking.
Nevertheless, training such a reranker is a very challenging task. As previously mentioned, it is unreasonable to use LLM’s feedback on each individual demonstration for training a reranker designed for $k$-shot selection, because demonstrations can influence each other. Additionally, constructing the ground truth ranking of a reranker tailored for $k$-shot selection requires finding the optimal $k$-shot permutation from the retrieved $n$ demonstrations. Theoretically, this requires using LLM to score total $\frac{n!}{(n-k)!}$ demonstration permutations, which is highly time-consuming and impractical. To overcome these challenges, we propose to construct a kind of dependency-aware training samples (a list of demonstrations with ranking labels) for reranker training. Specifically, given a retrieved demonstration set, we greedily select demonstrations from the set and annotate them with different ranking labels (from highest to lowest). Each time, the demonstration that maximizes the LLM’s feedback when concatenated with the already selected ones is chosen. This process not only considers the dependencies between current demonstration and previously selected ones, but also greatly reduces the number of LLM inferences.
To this end, we propose DemoRank, a Demonstration selection framework for passage Ranking, using a two-stage “retrieve-then-rerank” strategy. In this framework, we first train a demonstration retriever DRetriever based on LLM’s feedback for the ranking task. Then, we introduce a dependency-aware demonstration reranker DReranker to rerank the retrieved demonstrations. To address the challenges of its training, we propose a method to construct dependency-aware training samples that not only incorporates demonstration dependency but is also time-efficient.
Experiments on a series of ranking datasets prove the effectiveness of DemoRank, especially in few-shot ICL. Further analysis also demonstrates the contribution of each proposed component and DemoRank’s strong ability under different scenarios, including limited training data, different demonstration numbers, unseen datasets, etc.
The main contributions of our paper are summarized as follows: (1) To the best of our knowledge, we are the first to comprehensively discuss effective demonstration selection in passage ranking and propose DemoRank framework. (2) We propose a novel dependency-aware demonstration reranker and design a rational and efficient method for constructing its training data. (3) Besides in-domain performance, further experiments also demonstrate DemoRank’s generalization on unseen datasets.
2 Related Work
2.1 LLM for Passage Ranking
With the development of large language models (LLMs) in information retrieval Zhu et al. (2023), there have been many studies exploring how to utilize LLMs for the passage ranking task. In general, these studies can be divided into three categories: pointwise Liang et al. (2022); Sachan et al. (2022), pairwise Qin et al. (2023), and listwise methods Sun et al. (2023); Ma et al. (2023). Pointwise methods assess the relevance between a query and a single passage. A typical approach is relevance generation Liang et al. (2022); Zhuang et al. (2023a), which provides LLM with a query-passage pair and instructs it to output “Yes” if the passage is relevant to the query or “No” if not. The relevance score can be calculated based on the generation probability of the token “Yes”. Another approach of pointwise methods is query generation Sachan et al. (2022); Zhuang et al. (2023b), which calculates relevance score based on the log-likelihood of generating the query based on the passage. Pairwise methods compare two passages at a time and determine their relative relevance to a query, and listwise methods directly rank a passage list.
Despite promising results, these studies only focus on the zero-shot scenarios, with less emphasis on how to select effective demonstrations in few-shot scenarios. Manually written or rule-based selection Drozdov et al. (2023) is inflexible for ranking tasks. In this paper, we explore more effective demonstration selection approaches for ranking tasks. Previous studies Zhu et al. (2024) have revealed that relevance generation of the pointwise method is the most suitable method for passage ranking on open-source LLMs compared with other methods. Thus, we intend to use the relevance generation approach for passage ranking in this paper.
2.2 Demonstration Retrieval
A widely used demonstration selection approach is demonstration retrieval. Prior studies have explored using different retrievers for demonstration retrieval, which can be divided into two categories. One is utilizing off-the-shelf retrievers such as BM25 Agrawal et al. (2023) or dense retriever Liu et al. (2022). The other is to train a demonstration retriever using task-specific signals. For example, Rubin et al. (2022) propose to distill the LLM’s feedback to a dense retriever EPR for the semantic parsing task. Li et al. (2023) and Wang et al. (2023) propose to train the retriever iteratively on various NLP tasks. However, a common issue with these methods is that they directly choose the top-retrieved demonstrations, which may include redundant information and contribute little to the LLM’s understanding of relevance. In this paper, we take the demonstration dependencies into account and introduce a framework that first retrieves a list of demonstrations and then reranks in a dependency-aware manner, better aligning with the few-shot ICL in the ranking task.
3 Preliminaries
3.1 Relevance Generation for Ranking Task
Passage ranking aims to rank a list of retrieved passages based on their relevance to a query. Formally, given a query $q$ and a passage list $\left[p_{1},\dots,p_{n}\right]$, our task is to compute a relevance score $S(q,p_{i})$ for each passage. In the LLM-based relevance generation methods Liang et al. (2022); Zhuang et al. (2023a), an LLM is provided with a prompt consisting of a query and a passage, and instructed to output a binary label “Yes” or “No” to indicate whether the passage is relevant to the query or not. Then a softmax function is applied to the logits of tokens “Yes” and “No”, and the probability of the token “Yes” is used as the relevance score:
$Rs(q,p_{i})=\text{Pr}(\text{``Yes''}|T,q,p_{i}),$ (1)
where $T$ is the task description. Finally, the passages are ranked based on the relevance score $S(q,p_{i})$ in descending order.
3.2 In-context Learning in Ranking Task
In-context learning is a technique that inserts a few demonstrations into the prompt to help LLMs perform a task without updating parameters. In relevance generation task, given $k$ in-context demonstrations $\{z_{i}\}_{i=1}^{k}$, where $z_{i}=(\hat{q},\hat{p},\hat{y})$ is a triple consisting of a query, a passage and a binary output (“Yes” or “No”) indicating the relevance label, the relevance score $Rs(q,p_{i})$ could be calculated by:
$Rs(q,p_{i})=\text{Pr}(\text{``Yes''}|T,\{z_{i}\}_{i=1}^{k},q,p_{i}),$ (2)
where $T$ is the task description, which is used in ICL to help LLMs understand the task Zhu et al. (2024); Li et al. (2023).
Figure 2: An overview of our proposed framework DemoRank. DemoRank comprises two main components: DRetriever and DReranker. We train the DRetriever using demonstration candidates scored by LLM and construct a kind of dependency-aware training samples to train the DReranker. During inference, a retrieve-then-rerank pipeline is performed and the top-$k$ reranked ones are used for ICL.
4 The DemoRank Framework
As shown in Figure 2, our DemoRank framework follows a process of demonstration retrieval followed by dependency-aware reranking. The demonstration retriever DRetriever is trained using the demonstration candidates scored by LLM and the demonstration reranker DReranker is trained based on our constructed dependency-aware training samples. In this section, we elaborate on our demonstration pool construction, the pipeline of training, and inference.
4.1 Demonstration Pool Construction.
Given a passage ranking dataset (e.g., MS MARCO Nguyen et al. (2016)), we use its training set to construct our demonstration pool $\mathcal{P}$. For each query in the training set, we construct positive and negative demonstrations by pairing the query with its relevant and irrelevant passages respectively. To maintain the output label balance in the demonstration pool $\mathcal{P}$, the number of negative demonstrations of each query is set equal to its positive demonstrations.
4.2 Demonstration Retriever DRetriever
In this part, we train DRetriever to retrieve potentially useful demonstrations for subsequent demonstration reranking. We apply an LLM to score a set of demonstration candidates to obtain supervised signals and use them to train the retriever through a multi-task learning strategy.
Scoring with LLM
For a training input $I=(q,p)$ which contains a query-passage pair, we select a set of demonstrations from demonstration pool $\mathcal{P}$ as training candidates. Following previous studies Wang et al. (2023), we employ the BM25 algorithm to retrieve top-$b$ demonstrations. Due to the complex nature of passage ranking, the utility of a demonstration is not directly related to its similarity to the input Drozdov et al. (2023). To include more potential useful demonstrations for training, we also randomly sample another $B$ demonstrations from $\mathcal{P}$. The total number of training candidates is annotated as $N$ ($N=2*b$).
After that, we apply a frozen LLM scorer to score each demonstration $z_{i}$ for the training input $I$ using the following equation:
$f(z_{i},I)=\frac{\text{Pr}(y|T,z_{i},I)}{\sum_{y^{\prime}\in Y}\text{Pr}(y^{\prime}|T,z_{i},I)},$ (3)
where $y$ is the relevance label for the query-passage pair in $I$, $Y=\{``\text{Yes}",``\text{No}"\}$ is the label space and $T$ is the task description. In this paper, the LLM scorer uses the same model as the LLM passage ranker. Nevertheless, we also explored the transferability of LLM scorer on different LLM passage rankers (see Appendix C).
Training
Our DRetriever is based on bi-encoder architecture. Given the current training input $I=(q,p)$ and a candidate $z_{i}$, we use encoder $E_{I}$ and demonstration encoder $E_{z}$ to encode them respectively and calculate the similarity score as:
$S(I,z_{i})=E_{I}(I)^{\top}E_{z}(z_{i}),$ (4)
where the two encoders $E_{I}$ and $E_{z}$ share parameters and encode with average pooling.
Then we apply a contrastive loss $L_{\text{c}}$ to maximize the score between the training input $I$ and positive demonstration $z^{+}$ and minimize it for negative demonstration $z_{i}^{-}$. Here $z^{+}$ is the demonstration with the highest LLM score and $z_{i}^{-}$ are the remaining ones. The contrastive loss $L_{\text{c}}$ is calculated as:
$L_{\text{c}}=-\log\frac{e^{S(I,z^{+})}}{\sum_{z^{\prime}\in Z}e^{S(I,z^{\prime})}},$ (5)
where $Z=\{z^{+},z_{1}^{-},\dots,z_{N-1}^{-}\}$. Here we choose not to use in-batch negatives. The reasons are discussed in appendix D.
To make use of the fine-grained supervision of LLM’s feedback, we also consider a ranking loss RankNet Burges et al. (2005) to inject the ranking signal of candidates into training:
$L_{\text{r}}=\sum_{i,j}^{|Z|}\mathbb{1}_{r_{i}<r_{j}}*\log(1+e^{S(I,z_{j})-S(I,z_{i})}),$ (6)
where $r_{i}$ is the rank of $z_{i}$ in $Z$ when sorted in descending order by the LLM score.
The final loss function $L$ is defined as the weighted sum of $L_{c}$ and $L_{r}$:
$L=\lambda L_{c}+L_{r},$ (7)
where $\lambda$ is a pre-defined hyper-parameter.
4.3 Demonstration Reranker DReranker
Previous studies Wang et al. (2023); Rubin et al. (2022); Li et al. (2023) mainly use the top-$k$ retrieved demonstrations for ICL which ignores the demonstration dependencies and could be sub-optimal for ranking tasks. To mitigate this issue, we formulate the selection of the optimal $k$-shot permutation from retrieved demonstrations into a demonstration reranking problem and construct a novel dependency-aware training samples in an efficient way for the reranker’s training.
Algorithm 1 Constructing dependency-aware training samples
0:  Training input $I$, maximum iteration $K$.
0:  Dependency-aware training samples $Y$.
1:  $Y\leftarrow\{\}$, selected demonstrations $S\leftarrow[]$.
2:  Retrieve top-$M$ demonstrations $Z^{\text{r}}$
3:  for $y=K$ to $1$ do
4:     // $y$ is the current ranking label.
5:     $z^{*}=\operatorname{arg}\max_{z_{j}\in Z^{\text{r}}\backslash S}f(\left[S,z_{j}\right],I)$, using Equation (3)
6:     $S\leftarrow\left[S,z^{*}\right]$, $Y\leftarrow Y\cup\{(z^{*},y)\}$
7:  end for
8:  for $z_{j}$ in $Z^{\text{r}}\backslash S$ do
9:     $Y\leftarrow Y\cup\{(z_{j},0)\}$
10:  end for
11:  return  $Y$
Constructing Dependency-aware Training Samples.
To align with the aim of our DReranker, we propose constructing a dependency-aware training samples for training. Specifically, given a training input $I$, we use our trained DRetriever to retrieve top-$M$ demonstrations $Z^{\text{r}}$ from the demonstration pool. Then, we iteratively select demonstrations from $Z^{\text{r}}$ and annotate each of them with a ranking label, as Figure 2 shows. In each iteration, we select, from the unselected demonstrations in $Z^{\text{r}}$, the one that maximizes the LLM’s feedback when concatenated with already selected ones. Once a demonstration is selected, we append it to the training samples. This process considers previous demonstration sequence when selecting the current demonstration and approximates the optimal $k$-shot demonstration permutation incrementally, which is time-efficient and aligns with the few-shot setting. Note that as the number of iterations increases, the computational cost of LLM inference also increases. Due to limited computational resources, we set a maximum iteration number $K$. After the $K$-th iteration is completed, we annotate a ranking label from $K$ to 1 to each demonstration in the training sample according to their selection order and annotate 0 to the unselected demonstrations in $Z^{\text{r}}$. Algorithm 1 shows this procedure.
Training
After constructing the dependency-aware training sample, we obtain a ranking label for each demonstration candidate in $Z^{\text{r}}$. We employ a cross-encoder model to train our DReranker. The model takes as input the concatenation of training input $I$ and one candidate $z_{i}$ with a “[SEP]” token and outputs a prediction score $s_{i}$ using the representation of “[CLS]” token. Then we apply the RankNet loss function to optimize the reranker model, similar to Equation (6):
$L_{\text{r}}=\sum_{i,j}^{|Z^{\text{r}}|}\mathbb{1}_{y_{i}>y_{j}}*\log(1+e^{s_{j}-s_{i}}),$ (8)
where $y_{i}$ represents the ranking label of $z_{i}$. Note that our DReranker only receives an input and a single demonstration, without including dependent demonstrations, which may not fully capture the dependency-aware ranking labels. Nonetheless, this design saves inference time, making our DReranker more efficient. We plan to explore architectures that can model multiple dependent demonstrations efficiently in the future.
4.4 Inference
During inference, we first encode the entire demonstration pool $\mathcal{P}$ using our trained DRetriever and build the index. Then, given a test input $I^{\text{test}}=(q^{\text{test}},p^{\text{test}}_{i})$, we retrieve top-$M$ demonstrations using DRetriever and rerank them using our trained DReranker. Finally, we choose top-$k$ reranked demonstrations as the in-context demonstrations and concatenate them with the test input to calculate the relevance score. We perform this process for all retrieved passages of $q^{\text{test}}$ and rank these passages based on their relevance scores.
5 Experiments
5.1 Setting
Datasets
In our experiments, we train and evaluate our DemoRank on diverse ranking datasets, including HotpotQA Yang et al. (2018), NQ Kwiatkowski et al. (2019), FEVER Thorne et al. (2018) and MS MARCO Nguyen et al. (2016). We use their training set to train our models respectively and evaluate the models on the corresponding test set (for MS MARCO, the evaluation is conducted on its development set as well as two in-domain datasets, TREC DL19 Craswell et al. (2020b) and TREC DL20 Craswell et al. (2020a)).
Implementation Details
We use FLAN-T5-XL Chung et al. (2022) as the frozen LLM for demonstration scoring and passage ranking unless otherwise specified. During the training stage, the number of demonstration candidates for retriever and reranker ($N$ and $M$ respectively) are both set as 50. And the maximum iteration number $K$ in Section 4.3 is set as 4. During training, we apply e5-base-v2 Wang et al. (2022) and DeBERTa-v3-base He et al. (2023) to initialize our demonstration retriever and reranker respectively. Following previous studies Sun et al. (2023); Zhuang et al. (2023a), we use the top-100 passages retrieved by BM25 as the passages to rerank. Due to the limited space, more implementation details on model training and inference are listed in Appendix A.
Baselines
We compare our demonstration selection method with a series of baselines:
$\bullet$ Random: We randomly sample demonstrations from the demonstration pool $\mathcal{P}$ for each test input.
$\bullet$ DBS Drozdov et al. (2023): DBS is a rule-based selection approach based on query generation in passage ranking. It selects the demonstrations which are the most difficult for the LLM to predict. In this paper, we implemented the algorithm based on the relevance generation approach. We define a score for each demonstration as the probability of the LLM generating the corresponding relevance label given a query and passage. The demonstrations with the lowest scores are applied.
$\bullet$ K-means: K-means is another static demonstration selection approach. This method clusters all the demonstrations in the pool into $k$ clusters and then selects $k$ demonstrations closest to each cluster center for ICL. We use the E5 Wang et al. (2022) model to obtain the demonstration embeddings for clustering.
$\bullet$ BM25 Robertson and Zaragoza (2009): BM25 is a widely-used sparse retriever. We apply BM25 to retrieve demonstrations that are most similar to the test query.
$\bullet$ SBERT Reimers and Gurevych (2019): We use Sentence-BERT as the off-the-shelf demonstration retriever following Rubin et al. (2022) [^1] The checkpoint is from https://huggingface.co/sentence-transformers/paraphrase-mpnet-base-v2.. We use SBERT to encode all the demonstrations in the pool and retrieve the most similar demonstrations.
$\bullet$ E5 Wang et al. (2022): E5 is another off-the-shelf dense retriever. Following Wang et al. (2023), we use the same retrieval method as SBERT based on e5-base-v2 checkpoint [^2] https://huggingface.co/intfloat/e5-base-v2.
Method HotpotQA NQ FEVER DL19 DL20 MS MARCO Avg
Initial Order 63.30 30.55 65.13 50.58 47.96 22.84 46.73
0-shot 60.65 48.62 38.92 66.13 65.57 33.24 52.19
1-shot Random 59.71 48.69 38.41 66.76 65.35 33.53 52.08
K-means 59.62 48.68 37.96 66.45 65.30 33.59 51.93
DBS 60.34 49.05 38.96 66.83 65.79 33.54 52.42
BM25 61.46 49.53 40.43 65.08 65.86 33.73 52.68
SBERT 58.41 49.49 36.25 66.63 64.18 33.98 51.49
E5 61.70 49.49 39.96 66.48 65.20 33.79 52.77
DemoRank 65.64 52.11 44.16 68.64 67.38 35.03 55.49
3-shot Random 59.42 48.61 38.61 66.57 64.84 33.70 51.96
K-means 59.27 48.71 38.33 66.30 66.22 33.73 52.09
DBS 60.15 48.62 39.00 66.40 65.21 33.61 52.17
BM25 63.18 49.78 40.19 66.08 65.85 34.03 53.19
SBERT 58.38 49.23 36.80 66.67 65.07 33.71 51.64
E5 63.42 49.60 39.71 66.40 65.33 34.07 53.09
DemoRank 66.39 52.52 46.90 68.28 67.66 35.12 56.15
Table 1: Main results (NDCG@10) on different datasets. The best results are marked in bold and the column Avg represents the average performance of all datasets. The Initial Order represents the order of the top-100 passages retrieved by BM25.
5.2 Main Results
We compare DemoRank with baselines in 1-shot and 3-shot ICL respectively. Note that although DemoRank mainly focuses on few-shot scenarios, it can also work in 1-shot ICL, so we provide the performance of the 1-shot ICL as a reference. Table 1 shows the main results of our experiments. From the results, we draw the following observations: (1) Our framework DemoRank outperforms all the baselines significantly across all datasets. For example, in 3-shot ICL, DemoRank outperforms the second-best model E5 on HotpotQA by 3 points, and the second-best model BM25 on FEVER by about 7 points. It shows the DemoRank’s powerful ability to select demonstrations. (2) When expanding from 1-shot to 3-shot, DemoRank shows greater improvement on Avg metric compared to other baselines, indicating that our DemoRank can better enhance the ICL performance in few-shot scenario. (3) The similarity-based demonstration selection baselines (e.g., E5) outperform Random, K-means, and DBS baselines, but still lags far behind DemoRank, which proves the effectiveness of task-specific finetuning based on LLM’s feedback.
Method NQ DL19 FEVER Avg
Ablation study
- w/o DReranker 51.69 68.44 44.40 54.84
- w/o DTS 52.09 67.12 46.64 55.28
DemoRank 52.52 68.28 46.90 55.90
Using E5 as demonstration retriever
E5 49.60 66.40 39.71 51.90
$\text{DemoRank}_{\text{E5}}$ 50.74 67.37 41.76 53.29
Table 2: Results (NDCG@10) of different variants.
5.3 Analysis
In this section, we discuss different variants of DemoRank, compare DemoRank with supervised models, evaluate its performance on different demonstration numbers, and generalization on unseen datasets.
5.3.1 Different Variants of DemoRank
To understand the effectiveness of each component in DemoRank, we further evaluate different variants of DemoRank. We conduct the experiments on DL19, NQ sssand FEVER with 3-shot ICL, shown in Table 2. First, we remove our demonstration reranker DRanker and only use demonstrations retrieved by our demonstration retriever DRetriever, denoted as “- w/o DRanker”. We can see that removing DRanker causes about 1 point drop, which indicates that the reranked demonstrations are more useful for ICL. Secondly, to further validate the effectiveness of our dependency-aware training samples DTS in few-shot ICL, we introduce another variant that score each retrieved demonstration in $Z^{r}$ independently based on LLM, denoted as “- w/o DTS”. Without considering the demonstration dependency, this variant lags behind DemoRank by 0.62 points, which proves that the dependency-aware training samples align more with the few-shot ICL. Thirdly, we also replace our trained DRetriever with E5 in our framework to validate the training effectiveness of our DReranker on different demonstration retrievers, denoted as $\text{DemoRank}_{\text{E5}}$. From the results, we can see that $\text{DemoRank}_{\text{E5}}$ significantly improves E5, which proves that our DReranker’s training is flexible and not restricted by specific demonstration retriever. In addition, we also discuss the effectiveness of the ranking loss $L_{r}$ and in-batch negatives during DRetriever’s training in Appendix D.
Method Robust04 SCIDOCS DBPedia NEWS FiQA Quora NFCorpus Avg
Initial Order 40.70 14.90 31.80 39.52 23.61 78.86 33.75 37.59
monoBERT 44.18 15.99 41.70 44.62 32.06 74.65 34.97 41.17
0-shot 47.90 16.33 36.22 45.01 35.30 83.42 35.89 42.87
E5 46.49 16.78 37.72 45.40 35.38 84.13 35.44 43.05
DemoRank 48.14 16.90 39.76 46.47 35.93 83.96 36.14 43.90
Table 3: Results (NDCG@10) on BEIR. Best results are marked in bold. We use MS MARCO’s demonstration pool for retrieval and 3-shot ICL for E5 and DemoRank.
QNum Method MS MARCO DL19 DL20
0 0-shot 33.24 66.13 65.57
500K monoBERT 39.97 70.72 67.28
monoT5 40.05 70.58 67.33
DemoRank 35.12 68.28 67.66
20K monoBERT 30.69 63.61 59.32
monoT5 29.79 61.16 52.72
DemoRank 34.63 67.25 66.67
Table 4: Results (NDCG@10) on MS MARCO, DL19 and DL20. QNum represents the number of queries used in the MS MARCO training set.
5.3.2 Comparison with Supervised Reranker
The training of DemoRank is primarily based on queries in the training set, which can also be used to finetune a supervised model. In this part, we compare DemoRank with two supervised passage ranking models (monoBERT Nogueira and Cho (2019) and monoT5 Nogueira et al. (2020)) under different quantities of training queries. Training details of monoBERT and monoT5 are provided in Appendix B. We choose MS MARCO as the training set and NDCG@10 as the metric. We also report the 0-shot performance as a reference. The results are shown in Table 4. We can see that when provided with 500K queries, although DemoRank slightly outperforms monoBERT and monoT5 on DL20, it still lags behind them on DL19 and MS MARCO, indicating the advantages of supervised models when abundant training data is available. However, when the number of queries is limited to 20K, DemoRank significantly outperforms the two supervised models on three datasets and also shows a significant improvement over 0-shot baseline. This suggests that when training data is limited, DemoRank is more effective than supervised models, highlighting the potential of DemoRank in low-resource scenarios.
5.3.3 Different Demonstration Numbers
Demonstration number is often considered a key factor affecting ICL. In this part, we discuss the performance of our models under different demonstration numbers. We compare DemoRank with E5 baseline on FEVER and NQ datasets, using NDCG@10 as the metric. we also compare with our DRetriever to better understand the performance of our DReranker. The results are shown in Figure 3. We can see that both DRetriever and DemoRank outperform E5 consistently across different demonstration numbers, proving the effectiveness and robustness of our models. Besides, we can observe that as the demonstration number increases, the gap between DemoRank and DRetriever becomes more obvious (especially on FEVER), proving the effectiveness of dependency-aware demonstration reranking in few-shot ICL.
Figure 3: The impact of demonstration number.
5.3.4 Generalization on Unseen Datasets
One of the application scenarios of DemoRank is its generalization on unseen datasets. To prove this, we evaluate DemoRank trained on MS MARCO dataset on a series of BEIR datasets. We choose 0-shot, E5 demonstration retriever, and a supervised passage ranker MonoBERT Nogueira and Cho (2019), which is also trained on the MS MARCO dataset, for comparison. We use the demonstration pool from MS MARCO due to the lack of training sets in most BEIR datasets. As shown in Table 3, DemoRank outperforms the second-best model E5, by an average of about 1 point, proving its generalization ability. Furthermore, we also draw an interesting observation: despite using demonstrations from MS MARCO, DemoRank improves the 0-shot baseline across all datasets, indicating the potential of cross-dataset demonstrations in ICL.
6 Conclusion
In this paper, we explore how to select demonstrations for passage ranking task and propose DemoRank. We first trains a demonstration retriever with multi-task learning based on LLM’s feedback. Then, an reasonable and efficient method is propose to construct dependency-aware training samples, serving as the training data of the demonstration reranker. Experiments on various ranking datasets prove the effectiveness of DemoRank. Further analysis shows the effectiveness of each proposed component, the advantages compared to supervised models, and generalization on BEIR, etc.
Limitations
In this paper, we introduce a novel demonstration selection framework DemoRank for passage ranking task. We acknowledge several limitations in this paper that present opportunities for future work. First, due to limited computational resources, we can not conduct experiments with larger LLMs, such as those with 30B or even 70B parameters. Second, our framework is limited to pointwise passage ranking and lacks discussion on how demonstrations can be selected in pairwise and listwise passage ranking, which can be a promising direction to explore.
Appendix A Implementation Details of DemoRank
The hyper-parameters for training the demonstration retriever and reranker are shown in Table 5. For the construction of training input, we pair each query with one relevant passage and one irrelevant passage respectively, thus generating two training inputs. The passages labeled with 1 in the training set are used as relevant passages and the irrelevant ones are sampled from the top-100 passages retrieved by BM25. The number of queries used in each dataset is listed in Table 6. The maximum length of the queries and passages is set to 100 and 64, respectively.
During inference, for each test query-passage pair, we first use our DRetriever to retrieve top-50 demonstrations and then rerank them using our DReranker. The top-ranked demonstrations are used for ICL. The prompt we used consists of the instruction, demonstrations (one or more), and test input. For zero-shot, no demonstrations are included. The instructions and demonstrations we used are listed in Table 7 and Table 8 respectively. The instructions are used only for each test query-passage pair and the LLM scoring process. The test inputs have the same format as the demonstration.
Appendix B Training Details of Supervised Models
For fair comparison with DemoRank, we construct the training data by pairing each query with one relevant passage and one irrelevant passage respectively. As for monoBERT Nogueira and Cho (2019), we start training from a bert-large-uncased model and use a binary classification loss to optimize the model. As for monoT5 Nogueira et al. (2020), we initialize the model with T5-base model and finetune the model using generative loss. The training parameters are the same as the original paper.
Retriever Model Reranker Model
Initialization e5-base-v2 DeBERTa-v3-base
Optimizer AdamW AdamW
Learning Rate 3e-5 1e-5
Batch Size 8 8
Warmup Steps 400 400
Train Epochs 2 2
$\lambda$ 0.2 -
Table 5: Hyperparameters for training the demonstration retriever and reranker model.
      Query Number
      FEVER       150K
      NQ       150K
      HotpotQA       150K
      MS MARCO       200K
Table 6: The number of training queries for our framework DemoRank.
Dataset Instruction
FEVER Given an article and a claim, predict whether the article is relevant to the claim by outputting either Yes or No. If the article is relevant to the claim, output Yes; otherwise, output No.
NQ Given a passage and a question, predict whether the passage is relevant to the question by outputting either Yes or No. If the passage is relevant to the question, output Yes; otherwise, output No.
HotpotQA Given a passage and a question, predict whether the passage is relevant to the question by outputting either Yes or No. If the passage is relevant to the question, output Yes; otherwise, output No.
TREC DL19 Given a passage and a query, predict whether the passage is relevant to the query by outputting either Yes or No. If the passage is relevant to the query, output Yes; otherwise, output No.
TREC DL20 Given a passage and a query, predict whether the passage is relevant to the query by outputting either Yes or No. If the passage is relevant to the query, output Yes; otherwise, output No.
MS MARCO Given a passage and a query, predict whether the passage is relevant to the query by outputting either Yes or No. If the passage is relevant to the query, output Yes; otherwise, output No.
Table 7: The instructions used for different datasets.
Dataset Demonstration Format
[…]
Table 8: The demonstration format used for different datasets.
Appendix C Transferability across different LLM Ranker
In previous experiments, we used the same LLM (Flan-T5-XL) as the demonstration scorer and passage ranker. It is unknown whether the passage ranker could be replaced with other LLMs in the inference stage. In this section, we evaluate DemoRank’s transferability across different LLM rankers on several datasets and compare with several baselines, including 0-shot, Random, K-means, BM25, and E5. We experiment with Flan-T5-XXL [^3] https://huggingface.co/google/flan-t5-xxl (larger model size) and Llama-3-8B-Instruct [^4] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct (different model architecture) and the results are shown in Table 9. From the results, we can draw the following observations: (1) DemoRank outperforms all the baselines on Avg metric when using both Flan-T5-XXL and Llama-3-8B-Instruct as the LLM rankers, proving its strong transferability across different LLM rankers. (2) We observe that when using Flan-T5-XXL as LLM Ranker, DemoRank yields higher performance on FEVER, DL19, and MS MARCO (49.56, 68.74 and 35.90 respectively), compared with Flan-T5-XL (46.90, 68.28 and 35.12 respectively in Table 1). This shows DemoRank’s potential ability to improve passage ranking with larger-scale LLM rankers. (3) Comparing the overall 0-shot performance between Flan-T5-XL (see Table 1), Flan-T5-XXL and Llama-3-8B-Instruct, it is obvious that FlanT5 models perform better on average. This indicates that FlanT5 models are more suitable for passage ranking tasks, similar to findings from previous research Zhuang et al. (2023b).
HotpotQA NQ FEVER DL19 DL20 MS MARCO Avg
Initial Order 63.30 30.55 65.13 50.58 47.96 22.84 46.73
Flan-T5-XXL
0-shot 56.64 47.61 37.38 66.22 64.30 34.29 51.07
Random 58.31 48.51 39.56 67.47 65.31 35.15 52.39
K-means 58.75 48.86 39.37 67.40 65.47 35.24 52.52
BM25 60.66 50.47 43.89 66.82 65.67 35.15 53.78
E5 60.74 50.14 43.86 66.45 65.44 34.84 53.58
DemoRank 62.25 51.68 49.56 68.74 65.90 35.90 55.67
Llama-3-8B-Instruct
0-shot 55.93 36.24 27.53 58.47 55.10 28.09 43.56
Random 49.45 36.34 29.03 59.18 56.26 28.11 43.06
K-means 57.01 35.28 34.73 57.63 53.30 26.51 44.08
BM25 60.18 35.31 29.22 59.24 57.28 26.71 44.66
E5 60.09 36.76 28.64 57.36 53.09 26.98 43.82
DemoRank 60.89 35.47 45.24 60.36 56.45 28.54 47.83
Table 9: Results (NDCG@10) of different LLM ranker.
Appendix D Discussion on DRetriever’s Training
In this part, we conducted experiments to verify the rationale of Dretriever’s training. Firstly, we remove the ranking loss $L_{r}$ (Equation (6)) from training (denoted as “- w/o $L_{r}$”) and find a significant performance degradation on the three datasets. This indicates that the ranking signal in demonstration candidates is useful for demonstration retriever training. Besides, as we mentioned in Section 4.2, we do not apply in-batch negatives when calculating the contrastive loss $L_{c}$, which is different from previous studies Wang et al. (2023); Li et al. (2023); Karpukhin et al. (2020). To verify its rationale, we incorporate the in-batch negatives into the calculation of contrastive loss, denoted as “- w/ IBN”. From the results, we can see that the in-batch negatives do not bring significant improvement and even harm the retriever’s performance on DL19. This is because the utility of demonstrations in ranking tasks is not directly related to their similarity with the training input and the randomly sampled in-batch demonstrations may still contain valuable information and act as positive candidates, which is different from the assumption in passage retrieval Karpukhin et al. (2020). Thus, directly using in-batch negatives may introduce additional noise into the training process.
Method NQ DL19 FEVER Avg
- w/o $L_{r}$ 50.60 67.65 43.65 53.97
- w/ IBN 51.68 67.14 44.43 54.42
DRetriever 51.69 68.44 44.40 54.84
Table 10: The results (NDCG@10) of different training variants of DRetriever. We apply 3-shot ICL for each model.